Solve: What is the volume of a container that can hold 0.50 mole of gas at 25.0°C and 1.25atm?

Given: V1 = ?

n = 0.50 mole

T = 250C + 273.15= 298.15 K

P = 1.25 atm

R = 0.0821 L atm/ mol K


1
Expert's answer
2022-05-02T17:05:05-0400

Given:

P = 1.25 atm

V = unknown

n = 0.50 mol

R = 0.0821 L atm mol−1 K−1

T = 25.0°C + 273.15 = 298.15 K


Formula: PV = nRT


Solution:

The ideal gas law relates the amount of gas present to its pressure, volume, and temperature.

The ideal gas equation can be expressed as: PV = nRT

Rearrange the equation to solve for V:

V = nRT / P

Plug in the numbers:

V = (0.50 mol × 0.0821 L atm mol−1 K−1 × 298.15 K) / (1.25 atm) = 9.79 L

V = 9.79 L


Answer: The volume of a container is 9.79 L